OpenStudy (anonymous):

any ideas? Im clueless... Express cos(x + y) and sin(x + y) in terms of sin x, cos x, sin y, cos y. thanks

jimthompson5910 (jim_thompson5910):

hints: cos(A+B) = cos(A)cos(B) - sin(A)sin(B) sin(A+B) = sin(A)cos(B) + cos(A)sin(B)

OpenStudy (anonymous):

thanks jim_thompson5910 : )

jimthompson5910 (jim_thompson5910):

you're welcome